Please I need help with this problem:
1. The number of publishers in the usa has increased approximately linearly from 40,964 in 1993 to 73,200 publishers in 2003. Let n be the number of publishers at t years since 1990

(a) find an equation of a linear model to describe the data. (b) what is the slope? what does it mean in this situation? (c) what is the n intercept? what does it mean in this situation?

Answer by stanbon(26367)

Please I need help with this problem:
1. The number of publishers in the usa has increased approximately linearly from 40,964 in 1993 to 73,200 publishers in 2003. Let n be the number of publishers at t years since 1990
---
You have two points (t,n): (3,40,964) and (13,73,200)
slope = (73200-40964)/10 = 3223.6
intercept = ?
40,964 = 3223.6*3 + b
b = 31293.2
Equation: f(t) = 0.322.6n + 31293.2
---------------------------------------------------
The slope means the number of publishers increased by 3224 each year
after 1993
The n-intercept = (0,31293) means that in 1990 the number of
publishers was 31293 according to the model.
---------------------------------------------------
(a) find an equation of a linear model to describe the data. (b) what is the slope? what does it mean in this situation? (c) what is the n intercept? what does it mean in this situation?
===================================================
Cheers,
Stan H.